# Malle’s conjecture and Brauer groups of stacks

### v2

@@ 10.7. Abelian groups
Let now $G$ be a tame finite étale abelian group scheme. For $G$ constant, counting by discriminant was obtained in [63, 90] (for $k=\mathbb{Q}$ and $k$ general, respectively), and by conductor in [64, 91] (for $k=\mathbb{Q}$ and $k$ general, respectively). Moreover Wood’s work [91] obtained asymptotics for a suitable class of fair height functions. Non-constant $G$ have been considered in [4, 5, 32].
@@ 10.7.1. Brauer group of $BG$
For these facts see [78, Lem. 1.1, 1.2]. We use the map from Lemma 6.2.
@@ 10.7.3. Discriminant
We have $a(\Delta)=1/|G|(1-Q^{-1}),b(D)=(|G[Q]|-1)/[k(\mu_{Q}):k]$. Moreover $\Delta$ is fair if and only if $G=G[Q]$, i.e. $G$ is an elementary abelian group.
@@
Firstly as Conjecture 9.6 predicts, we should sort $G$-extensions according to the Iitaka fibration $BG\to B(G/G[Q])$. By Lemma 2.13 the fibres over rational points are isomorphic to the classifying stack of an inner twist of $G[Q]$. However since $G$ is abelian, any such inner twist is trivial. The discriminant becomes fair when restricted to a fibre, and the rational points in each fibre can be counted using the harmonic analysis techniques from [44, 45]. Conjecture 9.6 then predicts that the sum over all the fibres is absolutely convergent. This can be shown using the dominated convergence argument from Step 1 of the proof of [53, Thm. 1.1]. This gives a formula which is in agreement with Conjecture 9.6.
10.7.4. Other fair height functions
For $G$ abelian, in [91, §2.1] Wood calls a height function $H$ fair if $\langle\mathcal{M}(H)\rangle\cap G[r]$ generates $G[r]$ for all $r\in\mathbb{N}$. This is strictly stronger than our condition that $H$ be fair, which states only that $\mathcal{M}(H)$ generates $G$. Wood’s condition turns out to have strong consequences for the Brauer group.
@@ 10.7.4. Other fair height functions
Assume that $\langle\mathcal{M}(H)\rangle\cap G[r]$ generates $G[r]$ for all $r\in\mathbb{N}$. Then $\operatorname{Br}_{\mathcal{M}(L)}BG=\operatorname{Br_{un}}BG$.
@@
We now use that $G$ is constant to see that $k(\gamma)=k(\mu_{r})$ where $r$ is the order of $\gamma$. The assumption of the lemma implies that $\gamma=\sum_{m\in\langle\mathcal{M}(H)\rangle\cap G[r]}a_{m}m$ for some $a_{m}\in\mathbb{Z}$. For all $m\in G(-1)[r]$ we have $k(m)\subseteq k(\mu_{r})$.
As $\psi\in\operatorname{Br}_{\mathcal{M}(L),e}BG$ we have $\psi(\sigma)(m)=0$ for all $m\in\langle\mathcal{M}(H)\rangle(-1)\cap G(-1)[r]$ and all $\sigma\in\Gamma_{k(\mu_{r})}\subseteq\Gamma_{k(m)}$. As $\psi(\sigma)\in\operatorname{Hom}(G(-1)(k^{\mathrm{sep}}),\mathbb{Q}/\mathbb{Z})$ we have $\psi(\sigma)(\gamma)=\sum_{m\in\langle\mathcal{M}(H)\rangle\cap G[r]}a_{m}\psi(\sigma)(m)=0$ for all $\sigma\in\Gamma_{k(\mu_{r})}$. This implies that $\psi$ is unramified at $\gamma$, and we are done as $\gamma$ was arbitrary. ∎
Lemma 10.26 explains the fact that Wood was able in [91, Thm. 3.1] to state and prove a much simpler version of the equidistribution conjecture (Conjecture 9.10). Namely her conditions imply that she lay in case (2) of Proposition 9.11 and Remark 8.25. For general fair heights the equidistribution property is more complicated as the corresponding Brauer group can be larger; the conjecture in this case will be proven in forthcoming work of Tavernier [86].
An example of a height function which is fair in our sense, but not in the sense of Wood, would be a height on $B\mathbb{Z}/4\mathbb{Z}$ with weight function $1\mapsto 1,2\mapsto 2,3\mapsto 1$. This explicit example was provided by Wood and presented in Alberts’s paper [7, §7.6] as pathological to cast doubt on whether the definition of fair is correct, since the leading constant is a sum of two Euler products even though there is no Grunwald–Wang phenomenon. But the presence of this second Euler product is actually not pathological, and occurs as the relevant partially unramified Brauer group has two elements (modulo constants): the non-trivial element is given by $-4\in\operatorname{Br}_{1}B\mathbb{Z}/4\mathbb{Z}\cong H^{1}(k,\mu_{4})\cong k^{\times}/k^{\times 4}$ (this can be shown using Theorem 6.29(3)). This leads to a sum of Euler products as explained in Lemma 8.21.

### v3

@@ 10.7. Abelian groups
Let now $G$ be a tame finite étale abelian group scheme. For $G$ constant, counting by discriminant was obtained in [64, 92] (for $k=\mathbb{Q}$ and $k$ general, respectively), and by conductor in [65, 93] (for $k=\mathbb{Q}$ and $k$ general, respectively). Moreover Wood’s work [93] obtained asymptotics for a suitable class of balanced height functions. Non-constant $G$ have been considered in [4, 5, 32].
@@ 10.7.1. Brauer group of $BG$
For these facts see [80, Lem. 1.1, 1.2]. We use the map from Lemma 6.2.
@@ 10.7.3. Discriminant
We have $a(\Delta)=1/|G|(1-Q^{-1}),b(D)=(|G[Q]|-1)/[k(\mu_{Q}):k]$. Moreover $\Delta$ is balanced if and only if $G=G[Q]$, i.e. $G$ is an elementary abelian group.
@@
Firstly as Conjecture 9.6 predicts, we should sort $G$-extensions according to the Iitaka fibration $BG\to B(G/G[Q])$. By Lemma 2.13 the fibres over rational points are isomorphic to the classifying stack of an inner twist of $G[Q]$. However since $G$ is abelian, any such inner twist is trivial. The discriminant becomes balanced when restricted to a fibre, and the rational points in each fibre can be counted using the harmonic analysis techniques from [44, 45]. Conjecture 9.6 then predicts that the sum over all the fibres is absolutely convergent. This can be shown using the dominated convergence argument from Step 1 of the proof of [53, Thm. 1.1]. This gives a formula which is in agreement with Conjecture 9.6.
10.7.4. Other balanced height functions
For $G$ abelian, in [93, §2.1] Wood calls a height function $H$ fair if $\mathcal{M}(H)\cap G(-1)[r]$ generates $G[r]$ for all $r\in\mathbb{N}$. This is strictly stronger than our condition that $H$ be balanced, which states only that $\mathcal{M}(H)$ generates $G$. Wood’s condition turns out to have strong consequences for the Brauer group.
@@ 10.7.4. Other balanced height functions
Assume that $\mathcal{M}(H)\cap G(-1)[r]$ generates $G(-1)[r]$ for all $r\in\mathbb{N}$. Then $\operatorname{Br}_{\mathcal{M}(L)}BG=\operatorname{Br_{un}}BG$.
@@
We now use that $G$ is constant to see that $k(\gamma)=k(\mu_{r})$ where $r$ is the order of $\gamma$. The assumption of the lemma implies that $\gamma=\sum_{m\in\mathcal{M}(H)\cap G(-1)[r]}a_{m}m$ for some $a_{m}\in\mathbb{Z}$. For all $m\in G(-1)[r]$ we have $k(m)\subseteq k(\mu_{r})$.
As $\psi\in\operatorname{Br}_{\mathcal{M}(L),e}BG$ we have $\psi(\sigma)(m)=0$ for all $m\in\mathcal{M}(H)\cap G(-1)[r]$ and all $\sigma\in\Gamma_{k(\mu_{r})}\subseteq\Gamma_{k(m)}$. As $\psi(\sigma)\in\operatorname{Hom}(G(-1)(k^{\mathrm{sep}}),\mathbb{Q}/\mathbb{Z})$ we have $\psi(\sigma)(\gamma)=\sum_{m\in\mathcal{M}(H)\cap G(-1)[r]}a_{m}\psi(\sigma)(m)=0$ for all $\sigma\in\Gamma_{k(\mu_{r})}$. This implies that $\psi$ is unramified at $\gamma$, and we are done as $\gamma$ was arbitrary. ∎
Lemma 10.26 explains the fact that Wood was able in [93, Thm. 3.1] to state and prove a much simpler version of the equidistribution conjecture (Conjecture 9.10). Namely her conditions imply that she lay in case (2) of Proposition 9.11 and Remark 8.25. For general balanced heights the equidistribution property is more complicated as the corresponding Brauer group can be larger; the conjecture in this case will be proven in forthcoming work of Tavernier [88].
Example 10.27.
An example of a height function which is balanced, but not fair in the sense of Wood, would be a height on $B\mathbb{Z}/4\mathbb{Z}$ with weight function $1\mapsto 1,2\mapsto 2,3\mapsto 1$. This explicit example was provided by Wood and presented in Alberts’s paper [7, §7.6] as pathological to cast doubt on whether the definition of fair is correct, since the leading constant is a sum of two Euler products even though there is no Grunwald–Wang phenomenon. But the presence of this second Euler product occurs as the relevant partially unramified Brauer group has two elements (modulo constants): the non-trivial element is given by $-4\in\operatorname{Br}_{1}B\mathbb{Z}/4\mathbb{Z}\cong H^{1}(k,\mu_{4})\cong k^{\times}/k^{\times 4}$ (this can be shown using Theorem 6.29(3)). This leads to a sum of Euler products as explained in Lemma 8.21.
More details and further examples of related phenomena can be found in the work of Tavernier [88].